I'm really lost on this one when it comes to finding the roots, hopefully I'm on the right track!
Please factor the quadratic expression completely, and find the roots of the expression.
3. 135x2 - 222x + 91
(15x- 13)(9x-7)

Answer by venugopalramana(3286)

I'm really lost on this one when it comes to finding the roots, hopefully I'm on the right track!
Please factor the quadratic expression completely, and find the roots of the expression.
3. 135x2 - 222x + 91
135X^2-105X-117X+91=0
15X(9X-7)-13(9X-7)=0
(15x- 13)(9x-7)=0
15X-13=0......THAT IS .....15X=13....X=13/15
OR
9X-7=0........THAT IS ......9X=7.......X=7/9